Write the equation of the line passing through the point and perpendicular to the given line.

(5, 2), y=-1/2x +4
and
(6, 0), y=-2x+7

Answer by Nate(3500)

since the line in perpendicular you would take the opposite of the reciprical of the slope in y=-1/2x +4; the new slope would be 2
you need to use point slope form to find the equation:
y-(y value in ordered pair)=m(x-(x value in ordered pair))
y-2=2(x-5)
y-2=2x-10
y=2x-8 is your equations
to make long story short i will do the second one for you in my head
y=(1/2)x-3 should be your second equation